OpenStudy (anonymous):

Can Someone Help Me The Problem Says "The width of a rectangle is s. its length is 3s. How long is a diagonal of the rectangle? Express your answer in simplified radical form."

OpenStudy (wolf1728):

Using Pythagoren Theorem diagonal^2 = width^2 + length^2 diagonal^2 = s^2 + (3^2)*s^2 diagonal^2 = s^2 + 9s^2 diagonal^2 = 10s^2 diagonal = square root 10 * square root (s)

OpenStudy (anonymous):

woah

OpenStudy (wolf1728):

or square root of 10s

OpenStudy (anonymous):

ok thanks a lot